# Structure theorems for braided Hopf algebras

### v1

@@ 5.4 The topology of $\operatorname{BrPrim}$
[…]
@@
$1\to\{\ker_{Q}\}\to B_{n}\to\{Q\}\to 1$
[…]
@@
In principle, it is always possible to identify homology suboperads $H_{d(n-1)}(E_{2}^{\operatorname{fin}}(n))$ for a fixed integer $d$. However, since the homological dimension of $B_{n}$ is $n-1$, only $d=0$ or $1$ yield non-trivial operads.
@@ 5.5 The Woronowicz ideal and Nichols module
Assume that $\mathcal{I}\neq\mathfrak{W}$, so there is some $n$ and some element $\alpha\in\mathcal{I}(n)\setminus\mathfrak{W}(n)$. Then $\mathbf{S}_{n}(\alpha)\neq 0$, so by Proposition 1.12 there is some $V$ and $v_{1},\dots,v_{n}\in V$ with the property that $[\mathbf{S}_{n}(\alpha)](v_{1},\dots,v_{n})\neq 0$. Then $\alpha(v_{1},\dots,v_{n})\notin I_{n}=\mathcal{I}[V]_{n}$, which is a contradiction.
@@ 6.2 Profinitely braided bialgebras
The category $\operatorname{Alg}_{\widehat{\operatorname{BrAss}}}(\operatorname{pro}\!-\mathbf{C})$ becomes monoidal under $\operatorname{\widehat{\otimes}}$, with unit $\mathbf{1}$, and profinite completion defines a strictly monoidal functor
@@
The upper three categories are monoidal with respect to $\otimes$ or $\operatorname{\widehat{\otimes}}$, although $\operatorname{\widehat{\otimes}}$ does not obviously restrict to a monoidal product on $\operatorname{Alg}_{\widehat{\operatorname{BrAss}}}(\mathbf{C})$. Profinite completion and inclusion of constant objects are strictly monoidal functors (when this makes sense). In contrast, the forgetful functor $U$ is only lax monoidal, being a right adjoint.
@@ 6.3 Primitives in profinitely braided bialgebras
$[-,-]_{\sigma}:W\oplus V\to T(V\oplus W)\;\mbox{ given by }\;[w,v]_{\sigma}=w\otimes v-\sigma(w\otimes v).$
@@
[…]
@@ 6.4 Monoidal categories of $\mathcal{C}$-algebras
Note that the value $\mathcal{C}[0]$ of the Artin functor on 00 is
@@
Further, this is 00 if $\mathcal{C}$ is reduced, and $\mathbf{1}$ if $\mathcal{C}$ is unitary. Regardless of the value of $\mathcal{C}(0)$, there is an isomorphism $U_{\mathcal{C}}(\mathcal{C}[0])\cong\mathbf{1}$.
@@
We first show that $-\otimes_{\mathcal{C}}-$ is a bifunctor. Consider maps $f:L\to L^{\prime}$ and $g:M\to M^{\prime}$ of $\mathcal{C}$-algebras. There is a commuting diagram
@@
The existence of the map $f\otimes_{\mathcal{C}}g$ is implied by the fact that $U_{\mathcal{C}}(f)\operatorname{\widehat{\otimes}}U_{\mathcal{C}}(g)$ is a map of algebras (hence $\mathcal{C}$-algebras), and that it carries $L\operatorname{\widehat{\otimes}}1+1\operatorname{\widehat{\otimes}}M$ into $L^{\prime}\operatorname{\widehat{\otimes}}1+1\operatorname{\widehat{\otimes}}M^{\prime}$. It is easy to see from this construction that $-\otimes_{\mathcal{C}}-$ preserves composition and units.
@@
are restrictions of the unit isomorphisms in $(\operatorname{Alg}_{\widehat{\operatorname{BrAss}}}(\operatorname{pro}\!-\mathbf{C}),\operatorname{\widehat{\otimes}},\mathbf{1})$. Explicitly, note that $U_{\mathcal{C}}(\mathcal{C}[0])=\mathbf{1}$ is the unit for $\operatorname{\widehat{\otimes}}$, so
@@
is carried isomorphically onto $L$ (since $L$ is faithful) via the left unit isomorphism for $\operatorname{\widehat{\otimes}}$. An identical argument holds for the right unit.
The associativity isomorphism is also inherited from $\operatorname{\widehat{\otimes}}$. That is,
@@
Since the unit and associativity isomorphisms are inherited from $\operatorname{\widehat{\otimes}}$, they satisfy the required pentagonal and triangular coherence diagrams.
@@
which is induced (through the universal property of the enveloping algebra) by the inclusion of $\mathcal{C}$-algebras $L\otimes_{\mathcal{C}}M\subseteq U_{\mathcal{C}}(L)\operatorname{\widehat{\otimes}}U_{\mathcal{C}}(M)$. It is easy to verify the coherences required for monoidality, since we constructed $\otimes_{\mathcal{C}}$ in terms of $\operatorname{\widehat{\otimes}}$.
@@
$[-,-]_{\sigma}=\mu-\mu\circ\sigma:M\otimes L\to U_{\mathcal{C}}(L\otimes_{\mathcal{C}}M)$
@@ 6.5 Hopf $\mathcal{C}$-algebras
A Hopf $\mathcal{C}$-algebra in $\operatorname{pro}\!-\mathbf{C}$ is a coassociative, counital coalgebra [^15] The reader may complain that “Hopf” is inappropriate given that there is no analogue of an antipode in the definition. However, the term “$\mathcal{C}$-bialgebra” is also inappropriate, as it suggests that $L$ has operations and co-operations governed by $\mathcal{C}$. $L$ in the monoidal category of faithful $\mathcal{C}$-algebras in $\operatorname{pro}\!-\mathbf{C}$. A morphism of Hopf $\mathcal{C}$-algebras is a map $f:L\to M$ in $\operatorname{pro}\!-\mathbf{C}$ which is simultaneously a map of $\mathcal{C}$-algebras and of counital coalgebras.
@@
We note that if $\mathcal{C}$ is reduced, the target of the counit is the terminal object 00. In that case, there is a unique choice of $\epsilon$; then (6.14) is simply a constraint on $\delta$.
@@ 6.9 Minimal presentations of enveloping algebras
Let $\mathcal{I}\leq\widehat{\operatorname{BrAss}}$ be a left ideal, and let $\mathcal{D}:=\widehat{\operatorname{BrAss}}/\mathcal{I}$ be the associated quotient left module. Since $\mathcal{D}$ is not generally an operad unless $\mathcal{I}$ is also a right ideal, it does not have a well-defined category of algebras. Nonetheless, the value $\mathcal{D}[V]$ of the Artin functor associated to $\mathcal{D}$ defines a certain class of associative algebras. The purpose of this section is to give a presentation of sorts for these algebras, and more generally the enveloping algebra $U_{\mathcal{I}}(L)$ of an $\mathcal{I}$-algebra $L$ (which is $\mathcal{D}(L)$ if $L$ is abelian).
@@ 7.2 Operads with a perfect structure theory
We will say that a suboperad $\mathcal{C}\leq\widehat{\operatorname{BrAss}}$ has a perfect structure theory for $\mathbf{C}$ if, for all $V\in\operatorname{pro}\!-\mathbf{C}$, for all Hopf ideals $I\leq\widehat{T}(V)$, and all primitive elements $[x]\in\widehat{T}(V)/I$, $[x]$ has a representative $x$ which lies $\mathcal{C}[V]$.
@@ 8.2 Iteratively filtering by powers of primitives
[…]
@@
We induct on $k$; when $k=0$, the claim amounts to the statement that $P(A)\to P(A^{\operatorname{Pgr}})$ is an isomorphism in degree $1$. This holds for degree reasons: since $A^{\operatorname{Pgr}}$ is connected as a graded algebra (with respect to the second grading), the fact that the diagonal is graded implies that $A^{\operatorname{Pgr}}_{*,1}=P(A^{\operatorname{Pgr}})_{*,1}$. But $A^{\operatorname{Pgr}}_{*,1}$ vanishes if >1*>1∗ > 1, so this is concentrated in grading =1*=1∗ = 1: $A^{\operatorname{Pgr}}_{1,1}=P(A^{\operatorname{Pgr}})_{1,1}$. By definition, though, $A^{\operatorname{Pgr}}_{1,1}=P(A)_{1}$.
@@
Corollary 8.7 allows us to equip $A^{(\infty)}$ with a bigraded, profinitely braided bialgebra structure: to compute a product $xy$ of two bihomogenous elements $x\in A^{(\infty)}_{p,q}$ and $y\in A^{(\infty)}_{r,s}$, we may regard them as lying in $A^{(q+s)}$, compute their product there (which lies in $A^{(q+s)}_{p+r,q+s}$), and identify it with the corresponding element of $A^{(\infty)}_{p+r,q+s}$. A similar construction defines the diagonal. The verification of the axioms of a bialgebra may be done at finite stages.
@@ 9 The proof of Theorems 1 and 2
By assumption, there exists an object $V$, a Hopf ideal $I\leq\widehat{T}(V)$, and a class $[x]$ for which $[x]\in\widehat{T}(V)/I=:A$ is primitive, but for which no representative $x$ lies in $\mathcal{C}_{*}[V]$. Note that $\mathcal{C}_{*}[V]$ contains $V$, so in particular $x\notin V$.
@@
By definition $[x]\in\ker(a)$, so every representative $x\in\widehat{T}(V)$ must be in the kernel of the left composite. However, since no representative lies in $\mathcal{C}[V]$, the class of $x$ in $U_{\mathcal{C}_{*}}(P_{\mathcal{C}_{*}}(B))$ is nonzero, so must be nontrivial element of the kernel of the map to $B$.
@@ 10 A comparison with existing constructions
$[\dots]:P_{g_{1}}\otimes\cdots\otimes P_{g_{n}}\to P_{g_{1}+\cdots+g_{n}}$
[…]
@@
$[x_{1}\cdots x_{n}]=\sum_{\tau\in S_{n}}\left(\prod_{(i,j)}\zeta^{-1}\chi(g_{\tau(j)},g_{\tau(i)})\right)x_{\tau(1)}\cdots x_{\tau(n)}.$
@@
$[x_{1}\cdots x_{n}]=\sum_{\tau\in S_{n}}\zeta^{-\operatorname{inv}(\tau)}\mu_{n}(\widetilde{\tau}(x_{1}\otimes\cdots\otimes x_{n}))$
@@
More directly, Zhang-Zhang’s braided m-Lie algebras are the same as algebras over the braided Lie operad, $\operatorname{BrLie}$. If $L$ is a $\operatorname{BrLie}$-algebra, it is equipped with a binary operation as in (10.20) by definition. If we write $A=U_{\operatorname{BrLie}}(L)$ its enveloping algebra, then $ifL$ is faithful, the braided Lie bracket [,][-,-][ - , - ] is the restriction to $L$ of the operation
@@
[…]

### v2

@@ 5.4 The topology of $\operatorname{BrPrim}$
Let $\widetilde{E}_{2}(n)$ be the universal cover of $E_{2}(n)$. Since the fundamental group of $E_{2}(n)$ is the $n\textsuperscript{th}$ pure braid group, this collection forms a braided sequence in the category of topological spaces. In Example 3.1 of [Fie], Fiedorowicz defines [^14] Actually, Fiedorowicz does this for the operad of little rectangles, which is equivalent to the operad of little disks. a braided operad structure on $\widetilde{E}_{2}$. The operad structure maps $\circ_{i}$ are lifts of the corresponding maps in $E_{2}$ to the universal cover.
@@
$1\to\{\ker_{Q}\}\to B_{n}\to\{Q\}\to 1.$
Since homology is a lax monoidal functor, the homology $H_{*}(E_{2}^{\operatorname{fin}},k)$ forms a profinitely braided operad of graded $k$-modules; its $n\textsuperscript{th}$ term is the inverse system $\{H_{*}(\widetilde{E}_{2}(n)_{\ker_{Q}},k)\}$. It is possible to identify suboperads of $H_{*}(E_{2}^{\operatorname{fin}})$ by restricting the allowed degrees.
@@
In principle, it is always possible to identify homology suboperads $H_{d(n-1)}(E_{2}^{\operatorname{fin}}(n))$ for a fixed integer $d$. However, since the homological dimension of $B_{n}$ is n1n-1, only $d=0$ or $1$ yield non-trivial operads.
@@ 5.5 The Woronowicz ideal and Nichols module
Assume that $\mathcal{I}\neq\mathfrak{W}$, so there is some $n$ and some element $\alpha\in\mathcal{I}(n)\setminus\mathfrak{W}(n)$. Then $\mathbf{S}_{n}(\alpha)\neq 0$, so by Proposition 1.12 there is some $V$ and $v_{1},\dots,v_{n}\in V$ with the property that $\mathbf{S}_{n}(\alpha)](v_{1},\dots,v_{n})\neq 0$. Then $\alpha(v_{1},\dots,v_{n})\notin I_{n}=\mathcal{I}[V]_{n}$, which is a contradiction.
@@ 6.2 Profinitely braided bialgebras
[…]
@@
${\operatorname{Alg}_{\operatorname{BrAss}}(\operatorname{pro}\!-\mathbf{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}^\scriptstyle{\widehat{\cdot}}$${\operatorname{Alg}_{\operatorname{BrAss}}(\mathbf{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}^\scriptstyle{\widehat{\cdot}}$${\operatorname{Alg}_{\widehat{\operatorname{BrAss}}}(\operatorname{pro}\!-\mathbf{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{U}$${\operatorname{Alg}_{\widehat{\operatorname{BrAss}}}(\mathbf{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{U}$ (6.13)
@@ 6.3 Primitives in profinitely braided bialgebras
$_{\sigma}:W\oplus V\to T(V\oplus W)\;\mbox{ given by }\;[w,v]_{\sigma}=w\otimes v-\sigma(w\otimes v).$
@@
${\widehat{T}(V)_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{S}_{p,q}}$${\widehat{T}(V)\operatorname{\widehat{\otimes}}\widehat{T}(V)_{(p,q)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Nic}(V)_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta_{p,q}}$${\operatorname{Nic}(V)\operatorname{\widehat{\otimes}}\operatorname{Nic}(V)_{(p,q)}}$
@@ 6.4 Monoidal categories of $\mathcal{C}$-algebras
Note that the value $\mathcal{C}[0]$ of the Artin functor on 0 is
@@
Further, this is 0 if $\mathcal{C}$ is reduced, and $\mathbf{1}$ if $\mathcal{C}$ is unitary. Regardless of the value of $\mathcal{C}(0)$, there is an isomorphism $U_{\mathcal{C}}(\mathcal{C}[0])\cong\mathbf{1}$.
@@
We first show that $\otimes_{\mathcal{C}}-$ is a bifunctor. Consider maps $f:L\to L^{\prime}$ and $g:M\to M^{\prime}$ of $\mathcal{C}$-algebras. There is a commuting diagram
@@
The existence of the map $f\otimes_{\mathcal{C}}g$ is implied by the fact that $U_{\mathcal{C}}(f)\operatorname{\widehat{\otimes}}U_{\mathcal{C}}(g)$ is a map of algebras (hence $\mathcal{C}$-algebras), and that it carries $L\operatorname{\widehat{\otimes}}1+1\operatorname{\widehat{\otimes}}M$ into $L^{\prime}\operatorname{\widehat{\otimes}}1+1\operatorname{\widehat{\otimes}}M^{\prime}$. It is easy to see from this construction that $\otimes_{\mathcal{C}}-$ preserves composition and units.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$_{\sigma}=\mu-\mu\circ\sigma:M\otimes L\to U_{\mathcal{C}}(L\otimes_{\mathcal{C}}M)$
@@ 6.5 Hopf $\mathcal{C}$-algebras
A Hopf $\mathcal{C}$-algebra in $\operatorname{pro}\!-\mathbf{C}$ is a coassociative, counital coalgebra [^15] The reader may complain that “Hopf” is inappropriate given that there is no analogue of an antipode in the definition. However, the term $\mathcal{C}$-bialgebra” is also inappropriate, as it suggests that $L$ has operations and co-operations governed by $\mathcal{C}$. $L$ in the monoidal category of faithful $\mathcal{C}$-algebras in $\operatorname{pro}\!-\mathbf{C}$. A morphism of Hopf $\mathcal{C}$-algebras is a map $f:L\to M$ in $\operatorname{pro}\!-\mathbf{C}$ which is simultaneously a map of $\mathcal{C}$-algebras and of counital coalgebras.
@@
We note that if $\mathcal{C}$ is reduced, the target of the counit is the terminal object 0. In that case, there is a unique choice of $\epsilon$; then (6.14) is simply a constraint on $\delta$.
@@ 6.9 Minimal presentations of enveloping algebras
Let $\mathcal{I}\leq\widehat{\operatorname{BrAss}}$ be a left ideal, and let $\mathcal{D}:=\widehat{\operatorname{BrAss}}$$\mathcal{I}$ be the associated quotient left module. Since $\mathcal{D}$ is not generally an operad unless $\mathcal{I}$ is also a right ideal, it does not have a well-defined category of algebras. Nonetheless, the value $\mathcal{D}[V]$ of the Artin functor associated to $\mathcal{D}$ defines a certain class of associative algebras. The purpose of this section is to give a presentation of sorts for these algebras, and more generally the enveloping algebra $U_{\mathcal{I}}(L)$ of an $\mathcal{I}$-algebra $L$ (which is $\mathcal{D}(L)$ if $L$ is abelian).
@@ 7.2 Operads with a perfect structure theory
We will say that a suboperad $\mathcal{C}\leq\widehat{\operatorname{BrAss}}$ has a perfect structure theory for $\mathbf{C}$ if, for all $V\in\operatorname{pro}\!-\mathbf{C}$, for all Hopf ideals $I\leq\widehat{T}(V)$, and all primitive elements $\in\widehat{T}(V)/I, [x][x]$ has a representative $x$ which lies $\mathcal{C}[V]$.
@@ 8.2 Iteratively filtering by powers of primitives
The $k\textsuperscript{th}$ stage, $A^{(k)}$ carries a $k$-fold grading from the iteration of this procedure; if the original algebra $A$ was graded, then $A^{(k)}$ in fact is a $(k+1)$-fold graded, profinitely braided bialgebra. We will suppress all of these gradings except the last (coming from word length with respect to $P(A^{(k-1)})$) and the original grading on $A$, if it had one.
@@
We induct on $k$; when $k=0$, the claim amounts to the statement that $P(A)\to P(A^{\operatorname{Pgr}})$ is an isomorphism in degree $1$. This holds for degree reasons: since $A^{\operatorname{Pgr}}$ is connected as a graded algebra (with respect to the second grading), the fact that the diagonal is graded implies that $A^{\operatorname{Pgr}}_{*,1}=P(A^{\operatorname{Pgr}})_{*,1}$. But $A^{\operatorname{Pgr}}_{*,1}$ vanishes if >1*>1, so this is concentrated in grading =1*=1: $A^{\operatorname{Pgr}}_{1,1}=P(A^{\operatorname{Pgr}})_{1,1}$. By definition, though, $A^{\operatorname{Pgr}}_{1,1}=P(A)_{1}$.
@@
[…]
@@ 9 The proof of Theorems 1 and 2
By assumption, there exists an object $V$, a Hopf ideal $I\leq\widehat{T}(V)$, and a class $[x]$ for which $x]\in\widehat{T}(V)/I=:A$ is primitive, but for which no representative $x$ lies in $\mathcal{C}_{*}[V]$. Note that $\mathcal{C}_{*}[V]$ contains $V$, so in particular $x\notin V$.
@@
By definition $x]\in\ker(a)$, so every representative $x\in\widehat{T}(V)$ must be in the kernel of the left composite. However, since no representative lies in $\mathcal{C}[V]$, the class of $x$ in $U_{\mathcal{C}_{*}}(P_{\mathcal{C}_{*}}(B))$ is nonzero, so must be nontrivial element of the kernel of the map to $B$.
@@ 10 A comparison with existing constructions
$\dots]:P_{g_{1}}\otimes\cdots\otimes P_{g_{n}}\to P_{g_{1}+\cdots+g_{n}}$
associated to every $n\textsuperscript{th}$ root of unity $\zeta$ and $\zeta$-family” $(g_{1},\cdots,g_{n})$. The latter are tuples with the property that for each $i\neq j$, $\chi(g_{i},g_{j})\chi(g_{j},g_{i})=\zeta^{2}$.
@@
$x_{1}\cdots x_{n}]=\sum_{\tau\in S_{n}}\left(\prod_{(i,j)}\zeta^{-1}\chi(g_{\tau(j)},g_{\tau(i)})\right)x_{\tau(1)}\cdots x_{\tau(n)}.$
@@
$x_{1}\cdots x_{n}]=\sum_{\tau\in S_{n}}\zeta^{-\operatorname{inv}(\tau)}\mu_{n}(\widetilde{\tau}(x_{1}\otimes\cdots\otimes x_{n}))$
@@
More directly, Zhang-Zhang’s braided m-Lie algebras are the same as algebras over the braided Lie operad, $\operatorname{BrLie}$. If $L$ is a $\operatorname{BrLie}$-algebra, it is equipped with a binary operation as in (10.20) by definition. If we write $A=U_{\operatorname{BrLie}}(L)$ its enveloping algebra, then $ifL$ is faithful, the braided Lie bracket [,][-,-] is the restriction to $L$ of the operation
@@
The main result of [Ard11a] (Theorem 6.9) is an isomorphism U[]=limU[$^{[\infty]}=\varinjlim U^{[n]}\to A$. Comparing to our Theorem 1 (and particularly the proof of Lemma 9.1, which is inspired by this construction), we conclude that Ardizzoni’s enveloping algebra $U^{[\infty]}$ must be isomorphic as a bialgebra to our $U_{\mathfrak{W}_{*}}(P_{\mathfrak{W}_{*}}(A))$. More directly, (10.22) shows that $A$ is the quotient of $T(P(A))$ by the kernel $K$ of the composite displayed in that equation. We have a commuting diagram:
${T(P(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T(i)}$${T(P(A))/K=U^{[\infty]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{T(i)}}\scriptstyle{\cong}$${A}$${T(P_{\mathfrak{W}_{*}}(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U_{\mathfrak{W}_{*}}(P_{\mathfrak{W}_{*}}(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$